OpenStudy (anonymous):

Write an equation for: (3,-4) m=1/2

OpenStudy (anonymous):

y+4=1/2(x-3) y=1/2x-3/2-4 y=1/2x-11/2

OpenStudy (anonymous):

Thanks. I'm confused as to how you got the 11/2?

OpenStudy (anonymous):

3/2 + 4 = 3/2 +8/2 =11/2